OpenStudy (anonymous):

You need to buy laundry detergent and your favorite brand comes in three different sizes. Size 1 costs $7 and is enough for 28 loads, size 2 costs $14.72 and is enough for 64 loads, and size 3 costs $24 and is enough for 100 loads. Which size is the best deal? A. They are all the same B. Size 1 C. Size 2 D.

OpenStudy (kendricklamar2014):

\[\frac{ 7 }{ 28 } = $0.25\] \[Size~ 1: 0.25 \times 28 = $7 (No ~ savings)\] \[Size ~ 2: $0.25 \times 64 = $16 (Savings: $1.28)\] \[Size~ 3: $0.25 \times 100 = $25 (Savings: $1)\]

OpenStudy (anonymous):

Thank you!! :)

OpenStudy (kendricklamar2014):

Your welcome